1. If I roll a 4 sided die followed by flipping a coin, what is the probability that I roll an even number followed by getting heads? **I know the coin is (1/2), but what about the die?

2. If the probability Mr. Hansen makes a three point shot is 10%. What is the probability that Mr. Hansen misses two, three point shots in a row?
Answer #1

1)since there are 2 even number in 4 sided die , therefore P(even number) =2/4=1/2

P( roll an even number followed by getting heads ) =P(heads)*P(even number)=(1/2)*(1/2)=1/4

2)

P(missing two three point shots in a row ) =P(miss first)*P(miss 2nd) =(1-0.1)*(1-0.1)=0.9*0.9 =0.81